# Mapple: A Domain-Specific Language for Mapping Distributed Heterogeneous Parallel Programs

### v1

@@ 3.3. Transformation Primitives
$b_{t}=\begin{cases}a_{t}\qquad&t<p\text{ or }p<t<q\\ a_{p}\mathbin{\bmod}m.\mathtt{size}[p]\qquad&t=p\\ \left\lfloor\frac{a_{p}}{m.\mathtt{size}[p]}\right\rfloor\qquad&t=q\\ a_{t-1}\qquad&t>q\\ \end{cases}$
@@
$b_{t}=\begin{cases}a_{q}\qquad&\quad t=p\\ a_{p}\qquad&\quad t=q\\ a_{t}\qquad&\quad t\neq p\land t\neq q\\ \end{cases}$
$m^{\prime}=m.\mathtt{slice}(i,low,high)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1}]$
$b_{t}=\begin{cases}a_{i}+low\qquad&\quad t=i\\ a_{t}\qquad&\quad t\neq i\\ \end{cases}$
@@
We define the semantics of each Mapple transformation primitive in Figure 6, with the exception of the decompose primitive, which is deferred to Section 4. Each transformation primitive is a function that takes a processor space $m$ as input and returns a transformed processor space $m^{\prime}$, following the mapping illustrated on the right-hand side of Figure 6. We now describe each transformation in detail.
Figure 7. Common distributions expressed in Mapple. The shaded region in the iteration space is mapped to the corresponding shaded processor. The transformation code reshapes the original $(2,2)$ processor space into the desired processor space, which is then used in the user-defined mapping function.
@@
The split transformation takes two arguments: a dimension index $i$ and a splitting factor $d$. Given a processor space $m$ of shape $(s_{0},\dots,s_{n-1})$, the operation $m^{\prime}=m.\mathtt{split}(i,d)$ produces a new processor space $m^{\prime}$ of shape $(s_{0},\dots,d,s_{i}/d,\dots,s_{n-1})$. This transformation is invertible, allowing mappers to operate on $m^{\prime}$ while Mapple maps back to the original space $m$. The index relationship is:
$m^{\prime}[a_{0},\dots,a_{i},a_{i+1},\dots,a_{n}]=m[a_{0},\dots,a_{i}+a_{i+1}\times d,\dots,a_{n}].$
@@
The merge transformation takes two dimensions $p$ and $q$ of the processor space and fuses them into one. Given a processor space $m$ of shape $(s_{0},\dots,s_{n-1})$, the operation $m^{\prime}=m.\mathtt{merge}(p,q)$ produces a new processor space $m^{\prime}$ of shape $(s_{0},\dots,s_{p}\cdot s_{q},\dots,s_{n-1})$, where the two dimensions at positions $p$ and $q$ in $m$ are combined into a single dimension at position $p$ in $m^{\prime}$. This transformation is invertible. The index mapping is given by:
$m^{\prime}[a_{0},\dots,a_{n-2}]=m[a_{0},\dots,a_{p}\bmod s_{p},\dots,\lfloor a_{p}/s_{p}\rfloor,\dots,a_{n-2}].$
@@
Transformation primitives in Mapple can be composed sequentially. Consider a 2D processor space $m$, and let $m^{\prime}=m.\mathtt{split}(0,d)$, followed by $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$. The resulting processor space $m^{\prime\prime}$ is again 2D. We now derive the index transformation from $m^{\prime\prime}$ to the original space $m$ by sequentially applying the semantics of the merge and split transformations.
$m^{\prime\prime}[a_{0},a_{1}]=m^{\prime}[a_{0}\mathbin{\bmod}d,\lfloor a_{0}/d\rfloor,a_{1}]=m[(a_{0}\mathbin{\bmod}d)+\lfloor a_{0}/d\rfloor\times d,a_{1}].$
Because the expression $(a_{0}\mathbin{\bmod}d)+\lfloor a_{0}/d\rfloor\times d$ simplifies to $a_{0}$ for all integers, it follows that
$m^{\prime\prime}[a_{0},a_{1}]=m[a_{0},a_{1}].$
This demonstrates that the split and merge primitives are exact inverses when applied in sequence with the same parameters.
@@
The swap transformation exchanges two dimensions of the processor space. Combined with merge, which flattens two dimensions into one, it lets users control whether merging follows row-major or column-major order.
The slice transformation takes a dimension along with its lower and upper bounds, applying a constant offset to that dimension. It is useful for mapping an iteration space to a subset of the processor space, such as when assigning concurrent tasks to separate groups of processors.
@@
Using these primitives, Mapple can transform processor spaces in flexible ways. Figure 7 illustrates how Mapple supports common distribution patterns. Suppose we want to map a 2D iteration space onto a 2-node machine with 2 GPUs per node. In the figure, the shaded region of the iteration space is mapped to the corresponding shaded processor in the processor space.
For the block2D distribution, the mapping function is equivalent to the one in Figure 3, but simpler due to tuple arithmetic support. As seen in the mapping functions for block2D, block1D_x, and block1D_y, the only difference lies in the transformed processor space. Although all use 2D processor spaces, the dimensions differ, resulting in different block distributions. The same pattern applies to cyclic2D, cyclic1D_x, and cyclic1D_y.
@@ 4. Decompose Transformation Primitive
As discussed in Section 3.2.3, mismatches between the dimensionality of the iteration and processor spaces pose significant challenges for mapping decisions. In this section, we address this issue by analyzing its underlying principles and introducing a generalized transformation primitive, decompose. Section 4.1 analyzes a suboptimal heuristic used by a commonly used programming system to handle dimensionality mismatches. We then formally define the decompose primitive in Section 4.2, grounded in a communication volume analysis. Finally, Section 4.3 describes our search-based optimization algorithm and analyzes its complexity.
@@ 4.1. Suboptimal Existing Heuristics for Resolving Dimensionality Mismatch
We analyze how existing task-based systems handle dimensionality mismatches between iteration and processor spaces. Most frameworks, to the best of our knowledge, bypass this issue by linearizing both spaces and applying a default 1D block mapping. However, they rarely document the specifics of their linearization process, making direct comparisons challenging. An exception is Chapel (Chamberlain et al., 2007), a widely used parallel programming framework, which explicitly describes its approach to dimensionality mismatch. To reveal the limitations of this strategy, we present an intuitive example showing how it can lead to suboptimal mappings.
@@
Suppose we have 6 processors and a 2D iteration space. To match the dimensionality, we must split the 6 processors into a 2D tuple. There are four possible factorizations: $(6,1)$, $(3,2)$, $(2,3)$, and $(1,6)$. The existing algorithm used to determine the processor grid, as shown in Algorithm 1, does not consider the actual size of the iteration space.
The algorithm takes as input the number of processors to split, denoted by $d$, and the dimensionality of the iteration space, denoted by $k$. It returns an integer array of length $k$ whose product equals $d$. The algorithm follows a greedy strategy to compute this array, aiming to produce factors that are as balanced as possible in magnitude. It maintains a running product for each dimension and assigns the next prime factor of $d$ to the dimension with the smallest current product.
1
Input:
$d$ Number of processors
$k$ Dimensionality of the iteration space
2
3Function Greedy($d$, $k$):
5       factors $\leftarrow$ new int[$k$]
6       for $i\in\{0,\dots,k{-}1\}$ do
7             factors[$i$] $\leftarrow 1$
8      for $p\in$ primes do
9             $j\leftarrow$ ArgMin (factors) # Index of smallest current product
10             factors[$j$] $\leftarrow$ factors[$j$] ×$\times\ p$
11      Sort (factors, descending=True) # Sort for consistent ordering
12       return factors
Figure 8. For applications with spatial locality, the mapper computed by Algorithm 1 selects a fixed processor grid of $(3,2)$ for 2D block mapping. Data elements that must be transferred across processor boundaries are shown in orange. The $(12,18)$ iteration space incurs higher inter-processor communication volume than the $(18,12)$ iteration space, indicating that the chosen mapping is suboptimal for the former. A lower communication volume for the $(12,18)$ space could be achieved by using a $(2,3)$ processor grid instead.
@@ 11      Sort (factors, descending=True) # Sort for consistent ordering
In the example above, the algorithm selects the grid $(3,2)$ by prioritizing balanced factorization, without considering the shape of the iteration space. This can lead to suboptimal mappings. We illustrate this with two iteration spaces, $(12,18)$ and $(18,12)$, corresponding to an application with spatial locality, as shown in Figure 8. Each iteration space is mapped to the processor grid $(3,2)$ using the block2D function, which partitions the iteration space into rectangular blocks assigned to each processor. The orange regions in the figure indicate data that must be transferred across processor boundaries due to this mapping.
@@
For the $(12,18)$ iteration space, the total inter-processor communication volume is 96 elements, while for the $(18,12)$ iteration space it is 84 elements. This difference highlights the sensitivity of communication cost to how the iteration space is aligned with the processor grid. If the algorithm had taken the shape of the iteration space into account and chosen the processor grid $(2,3)$ for the $(12,18)$ space, the communication volume could have matched the more efficient 84-element case, avoiding unnecessary communication overhead.
@@ 4.2. Formal Definition of Decompose and Intuitive Solution
We formally define the optimization problem that the decompose primitive aims to solve, based on an analysis of inter-processor communication volume. We also present the underlying intuition to clarify how the solution addresses this problem.
@@
Conceptually, decompose is a shorthand for applying a sequence of split transformations. Specifically, applying $m_{k}=m_{1}.\texttt{decompose}(i,T)$ is equivalent to performing $m_{n+1}=m_{n}.\texttt{split}(i+n-1,d_{i_{n}})$ for $1\leq n<k$.
There can be multiple ways to factor $d_{i}$ into $k$ natural numbers $d_{i_{1}},\ldots,d_{i_{k}}$. The decompose primitive selects a factorization by solving the following optimization problem:
minimize $\sum_{m=1}^{k}\frac{d_{i_{m}}}{l_{m}}$
subject to $\prod_{m=1}^{k}d_{i_{m}}=d_{i},\quad d_{i_{m}}\in\mathbb{N}\;\;\forall m$
An equivalent formulation introduces the workload vector $w_{m}=\frac{l_{m}}{d_{i_{m}}}$ for $1\leq m\leq k$, where $w_{m}$ represents the amount of iteration space assigned to each processor in the $m$-th dimension. Rewriting the problem in terms of the workload vector gives:
minimize $\sum_{m=1}^{k}\frac{1}{w_{m}}$
subject to $\prod_{m=1}^{k}w_{m}=\frac{\prod_{m=1}^{k}l_{m}}{d_{i}},\quad\frac{l_{m}}{w_{m}}\in\mathbb{N}\;\;\forall m$
@@
We now illustrate the optimization problem using the two examples shown in Figure 9. The left-hand side depicts a 2D case with $k=2$ and processor count $d_{i}=4$. To compute the communication volume, we count the number of elements transferred across processor boundaries. This volume equals $2\cdot(w_{1}+w_{2})\cdot d_{i}-2\cdot(l_{1}+l_{2})$, where the first term is the total perimeter of the $d_{i}$ blocks of size $(w_{1},w_{2})$, and the second term is the perimeter of the entire iteration space of size $(l_{1},l_{2})$.
@@
On the right-hand side, the figure shows a 3D example with $k=3$ and $d_{i}=16$. Here, the communication volume corresponds to the total inter-block surface area. The surface area $S$ of the $d_{i}$ cuboids of size $(w_{1},w_{2},w_{3})$ is given by:
$2S=2(w_{1}w_{2}+w_{1}w_{3}+w_{2}w_{3})\cdot d_{i}-2(l_{1}l_{2}+l_{1}l_{3}+l_{2}l_{3}).$
Given fixed input dimensions $l_{1},l_{2},l_{3}$ and processor count $d_{i}$, the term $w_{1}w_{2}w_{3}=\frac{l_{1}l_{2}l_{3}}{d_{i}}$ is constant. Thus, minimizing $S$ is equivalent to minimizing $w_{1}w_{2}+w_{1}w_{3}+w_{2}w_{3}$. Under the fixed-product constraint, this can be transformed into minimizing $\frac{1}{w_{1}}+\frac{1}{w_{2}}+\frac{1}{w_{3}}$, again aligning with the objective of our optimization formulation in the 3D case.
This analysis generalizes to any $k$-dimensional case. Let $S$ denote the total inter-surface area in the $k$-dimensional setting. Then,
@@
This is the well-known arithmetic-geometric mean (AM-GM) inequality. It states that the arithmetic mean of a set of positive numbers is always greater than or equal to their geometric mean, with equality only when all values are equal.
Note that in the optimization problem, the term $\frac{\prod_{m=1}^{k}l_{m}}{d_{i}}$ is a constant determined by the input size and the number of processors. We can now apply the AM–GM inequality to the objective: $\frac{1}{k}\sum_{m=1}^{k}\frac{1}{w_{m}}\geq\left(\prod_{m=1}^{k}\frac{1}{w_{m}}\right)^{1/k}$. Using the constraint $\prod_{m=1}^{k}w_{m}=\frac{\prod_{m=1}^{k}l_{m}}{d_{i}}$, we obtain:
$\sum_{m=1}^{k}\frac{1}{w_{m}}\geq k\cdot\left(\frac{d_{i}}{\prod_{m=1}^{k}l_{m}}\right)^{1/k}.$
Equality is achieved when all terms $\frac{1}{w_{m}}$ are equal, which implies $w_{1}=w_{2}=\cdots=w_{k}$. In practice, however, exact equality may not be attainable due to the integrality constraint that each $\frac{l_{m}}{w_{m}}$ must be a natural number.
In the 2D iteration space example shown in Figure 8, the mapping for the $(18,12)$ space is optimal because the workload vector is $(w_{1},w_{2})=\left(\frac{l_{1}}{d_{1}},\frac{l_{2}}{d_{2}}\right)=(6,6)$, where the two components are equal. Similarly, the 3D example in Figure 9 uses an iteration space of $(4,8,4)$ mapped onto 16 processors. The solution shown in the figure yields a workload vector $(w_{1},w_{2},w_{3})=(2,2,2)$, which satisfies the sufficient condition for attaining the minimum communication volume.
These examples suggest that the optimal solution is achieved when the iteration space is divided among the processor dimensions in a balanced manner. However, in practice, this ideal may not be attainable due to the integrality constraint that each $\frac{l_{m}}{w_{m}}$ must be a natural number. To address this, we introduce a search-based algorithm for solving the optimization problem, which we describe in detail in Section 4.3.

### v2

@@ 3.3. Transformation Primitives
[…]
@@
[…]
@@
We define the semantics of each Mapple transformation primitive in Figure 6, with the exception of the decompose primitive, which is deferred to Section 4. Each transformation primitive is a function that takes a processor space $m$ as input and returns a transformed processor space $m^{\prime}$, following the mapping illustrated on the right-hand side of Figure 6. We now describe each transformation.
@@
The split transformation takes two arguments: a dimension index $i$ and a splitting factor $d$. Given a processor space $m$ of shape $(s_{0},\dots,s_{n-1})$, the operation $m^{\prime}=m.\mathtt{split}(i,d)$ produces a new processor space $m^{\prime}$ of shape $(s_{0},\dots,d,s_{i}/d,\dots,s_{n-1})$. Like other processor space transformations, this transformation is invertible; in our implementation, Mapple uses the inverse to translate from mapper operations on $m^{\prime}$ to the original space $m$.
@@
The merge transformation takes two dimensions $p$ and $q$ of the processor space and fuses them into one. Given a processor space $m$ of shape $(s_{0},\dots,s_{n-1})$, the operation $m^{\prime}=m.\mathtt{merge}(p,q)$ produces a new processor space $m^{\prime}$ of shape $(s_{0},\dots,s_{p}\cdot s_{q},\dots,s_{n-1})$, where the two dimensions at positions $p$ and $q$ in $m$ are combined into a single dimension at position $p$ in $m^{\prime}$. The index mapping is given by: $m^{\prime}[a_{0},\dots,a_{n-2}]=m[a_{0},\dots,a_{p}\bmod s_{p},\dots,\lfloor a_{p}/s_{p}\rfloor,\dots,a_{n-2}]$.
@@
Transformation primitives in Mapple can be composed sequentially. Consider a 2D processor space $m$, and let $m^{\prime}=m.\mathtt{split}(0,d)$, followed by $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$. The resulting processor space $m^{\prime\prime}$ is again 2D. We now derive the index transformation from $m^{\prime\prime}$ to the original space $m$ by sequentially applying the semantics of the merge and split transformations: $m^{\prime\prime}[a_{0},a_{1}]=m^{\prime}[a_{0}\mathbin{\bmod}d,\lfloor a_{0}/d\rfloor,a_{1}]=m[(a_{0}\mathbin{\bmod}d)+\lfloor a_{0}/d\rfloor\times d,a_{1}]$. Because the expression $(a_{0}\mathbin{\bmod}d)+\lfloor a_{0}/d\rfloor\times d$ simplifies to $a_{0}$ for all integers, it follows that $m^{\prime\prime}[a_{0},a_{1}]=m[a_{0},a_{1}]$merge is the inverse of split.
@@
The swap transformation exchanges two dimensions of the processor space. Combined with merge, which flattens two dimensions into one, users can control whether merging follows row-major or column-major order.
@@
Using these primitives, Mapple can transform processor spaces in flexible ways. Figure 7 illustrates how Mapple supports common distribution patterns. Suppose we want to map a 2D task space onto a 2-node machine with 2 GPUs per node. In the figure, the shaded region of the task space is mapped to the corresponding shaded processor in the processor space.
For the block2D distribution, the mapping function is equivalent to the one in Figure 3, but is simpler thanks to Mapple’s support for tuple arithmetic. Tuple operands must have the same shape. In Figure 7, both the transformed processor spaces and the task space are 2D. In particular, the * operator flattens tuples for indexing: for example, m[*(1,2)] is equivalent to m[1,2]. Mapple supports both elementwise arithmetic on individual tuple dimensions and operations over entire tuples, enabling users to express index mappings more concisely and intuitively.
Figure 7. Common distributions expressed in Mapple. The shaded region in the task space is mapped to the corresponding shaded processor. The transformation code reshapes the original $(2,2)$ processor space into the desired processor space, which is then used in the user-defined mapping function.
@@ 4. Decompose Transformation Primitive
As discussed in Section 3.2.3, mismatches between the dimensionality of the task and processor spaces pose significant challenges for mapping decisions. In this section, we address this issue by analyzing its underlying principles and introducing a generalized transformation primitive, decompose. Section 4.1 analyzes a suboptimal, but standard, heuristic to handle dimensionality mismatches. We then formally define the decompose primitive in Section 4.2, grounded in a communication volume analysis. Finally, Section 4.3 describes our search-based optimization algorithm and analyzes its complexity.
@@ 4.1. Suboptimal Existing Heuristics for Resolving Dimensionality Mismatch
We examine how existing task-based systems address dimensionality mismatches between task and processor spaces. Most frameworks sidestep this issue by linearizing both spaces and applying a default 1D block mapping. However, they rarely document their linearization procedures, making direct comparison difficult. An exception is Chapel (chamberlain2007parallelchapel), a widely used parallel programming framework, which explicitly describes its approach. To expose the limitations of this strategy, we present an example showing how it can lead to suboptimal mappings.
@@
Suppose we have 6 processors and a 2D task space. To match the dimensionality, we must split the 6 processors into a 2D tuple. There are four possible factorizations: $(6,1)$, $(3,2)$, $(2,3)$, and $(1,6)$. The existing heuristic used to determine the processor grid, as shown in Algorithm 1 of the supplementary material, does not consider the shape of the task space.
@@
The heuristic follows a greedy strategy, aiming to produce factors that are as balanced as possible in magnitude. In the example above, the algorithm selects the grid $(3,2)$ by prioritizing balanced factorization. Now consider two task spaces, $(12,18)$ and $(18,12)$, for an application with nearest neighbor locality (i.e., data is exchanged among immediate neighbors in task space), as shown in Figure 8. Each task space is mapped to the processor grid $(3,2)$ using the block2D function, which partitions the task space into rectangular blocks assigned to each processor. The orange regions in the figure indicate data that must be transferred across processor boundaries due to this mapping.
@@
For the $(12,18)$ task space, the total inter-processor communication volume is 96 elements, while for the $(18,12)$ task space it is 84 elements. This difference highlights the sensitivity of communication cost to how the task space is aligned with the processor grid. If the algorithm had taken the shape of the task space into account and chosen the processor grid $(2,3)$ for the $(12,18)$ space, the communication volume could have matched the more efficient 84-element case, avoiding unnecessary communication overhead.
Figure 8. For this application, the mapper from Algorithm 1 (see supplementary material) selects a fixed $(3,2)$ processor grid for 2D block mapping. Orange regions indicate inter-processor data transfers. The $(12,18)$ task space incurs higher communication than $(18,12)$, showing that the chosen grid is suboptimal; a $(2,3)$ grid would yield lower communication.
@@ 4.2. Formal Definition of Decompose and Intuitive Solution
We formally define the optimization problem that the decompose primitive aims to solve, based on an analysis of inter-processor communication volume. We also present the underlying intuition to clarify how the solution addresses this problem. Implicit in all solutions to this problem (including the greedy heuristic described above) are two assumptions: that tasks in the task space have uniform cost (i.e., are load balanced) and that there is locality in the communication pattern, so that partitioning the task space into compact regions is desirable to minimize communication.
@@
Conceptually, decompose is a shorthand for applying a sequence of split transformations. Specifically, applying $m_{k}=m_{1}.\texttt{decompose}(i,T)$ is equivalent to performing $m_{n+1}=m_{n}.\texttt{split}(i+n-1,d_{i_{n}})$ for $1\leq n<k$. There can be multiple ways to factor $d_{i}$ into $k$ natural numbers $d_{i_{1}},\ldots,d_{i_{k}}$. The decompose primitive selects a factorization by solving the following optimization problem:
$\min_{\{d_{i_{m}}\in\mathbb{N}\}}\sum_{m=1}^{k}\frac{d_{i_{m}}}{l_{m}}\quad\text{s.t.}\quad\prod_{m=1}^{k}d_{i_{m}}=d_{i}.$
An equivalent formulation introduces the workload vector $w_{m}=\frac{l_{m}}{d_{i_{m}}}$ for $1\leq m\leq k$, where $w_{m}$ represents the portion of the task space assigned to each processor in the $m$-th dimension. Rewriting the problem in terms of the workload vector gives:
$\min_{\{w_{m}\}}\sum_{m=1}^{k}\frac{1}{w_{m}}\quad\text{s.t.}\quad\prod_{m=1}^{k}w_{m}=\frac{\prod_{m=1}^{k}l_{m}}{d_{i}},\;\frac{l_{m}}{w_{m}}\in\mathbb{N}.$
@@
We now illustrate the optimization problem using the two examples shown in Figure 9. The left-hand side depicts a 2D case with $k=2$ and processor count $d_{i}=4$. To compute the communication volume, we count the number of elements transferred across processor boundaries. This volume equals $2\cdot(w_{1}+w_{2})\cdot d_{i}-2\cdot(l_{1}+l_{2})$, where the first term is the total perimeter of the $d_{i}$ blocks of size $(w_{1},w_{2})$, and the second term is the perimeter of the entire task space of size $(l_{1},l_{2})$.
@@
On the right-hand side, the figure shows a 3D example with $k=3$ and $d_{i}=16$. Here, the communication volume corresponds to the total inter-block surface area. The surface area $S$ of the $d_{i}$ cuboids of size $(w_{1},w_{2},w_{3})$ is given by $2S=2(w_{1}w_{2}+w_{1}w_{3}+w_{2}w_{3})\cdot d_{i}-2(l_{1}l_{2}+l_{1}l_{3}+l_{2}l_{3})$. Given fixed input dimensions $l_{1},l_{2},l_{3}$ and processor count $d_{i}$, the term $w_{1}w_{2}w_{3}=\frac{l_{1}l_{2}l_{3}}{d_{i}}$ is constant. Thus, minimizing $S$ is equivalent to minimizing $w_{1}w_{2}+w_{1}w_{3}+w_{2}w_{3}$. Under the fixed-product constraint, this can be transformed into minimizing $\frac{1}{w_{1}}+\frac{1}{w_{2}}+\frac{1}{w_{3}}$, again aligning with the objective of our optimization formulation in the 3D case. This analysis generalizes to any $k$-dimensional case. Let $S$ denote the total inter-surface area in the $k$-dimensional setting. Then,
@@
This is the well-known arithmetic-geometric mean (AM-GM) inequality. It states that the arithmetic mean of a set of positive numbers is always greater than or equal to their geometric mean, with equality only when all values are equal. In the optimization problem, the term $\frac{\prod_{m=1}^{k}l_{m}}{d_{i}}$ is a constant determined by the input size and the number of processors. We can now apply the inequality to the objective: $\frac{1}{k}\sum_{m=1}^{k}\frac{1}{w_{m}}\geq\left(\prod_{m=1}^{k}\frac{1}{w_{m}}\right)^{1/k}$. Given $\prod_{m=1}^{k}w_{m}=\frac{\prod_{m=1}^{k}l_{m}}{d_{i}}$, we obtain $\sum_{m=1}^{k}\frac{1}{w_{m}}\geq k\cdot\left(\frac{d_{i}}{\prod_{m=1}^{k}l_{m}}\right)^{1/k}$. Equality is achieved when all terms $\frac{1}{w_{m}}$ are equal, which implies $w_{1}=w_{2}=\cdots=w_{k}$. In practice, however, exact equality may not be attainable because each $\frac{l_{m}}{w_{m}}$ must be a natural number.
In the 2D task space example shown in Figure 8, the mapping for the $(18,12)$ space is optimal because the workload vector is $(w_{1},w_{2})=\left(\frac{l_{1}}{d_{1}},\frac{l_{2}}{d_{2}}\right)=(6,6)$, where the two components are equal. Similarly, the 3D example in Figure 9 uses a task space of $(4,8,4)$ mapped onto 16 processors. The solution shown in the figure yields a workload vector $(w_{1},w_{2},w_{3})=(2,2,2)$, which satisfies the sufficient condition for attaining the minimum communication volume.
These examples suggest that the optimal solution is achieved when the task space is divided among the processor dimensions in a balanced manner. However, in practice, this ideal may not be attainable due to the integrality constraint that each $\frac{l_{m}}{w_{m}}$ must be a natural number. We introduce a search-based algorithm for solving the optimization problem in Section 4.3.
As mentioned above, our analysis assumes locality in communication (e.g., such as nearest-neighbor communication), which is very common, but not universal, in practice. Other typical patterns, including anisotropic halo and all-to-all exchanges, are discussed in Section A.3 (see supplementary material). For algorithms without spatial locality or highly unbalanced tasks, decompose may be ineffective. In such cases, a good Mapple mapper may rely on other features, such as load-balancing (see Section 7).